Logan saves $5.50 in dimes and quarters over a week. He has 20 more dimes than quarters. Find the number of dimes and quarters he saves.
Answer by ReadingBoosters(3246)

x = dimes
y = quarters
.1x+.25y=5.50
x=20+y
.1(20+y) + .25y = 5.50
2 + .1y + .25y = 5.50
.35y = 3.50
y = 10 quarters
x = 20+10=30 dimes